OpenStudy (lgbasallote):

what is the inverse of f(x) = x^2 if the domain is all integers?

hartnn (hartnn):

range of f(x) cannot be all integers. so domain of inverse of f(x) can't be all integers

OpenStudy (lgbasallote):

....

OpenStudy (anonymous):

To find an inverse, you should exchange the y and x values in your equation for f abd solve for y\[y=x ^{2}\]\[x=y ^{2}\]Solve from there.

OpenStudy (anonymous):

hartnn, I think the domain refers to f of x 's domain...?

hartnn (hartnn):

if you allow complex values, then its \(\sqrt x\)

OpenStudy (lgbasallote):

enlighten me

hartnn (hartnn):

like square of complex numbers can be negative. then x^2 can take negative integers and \(\sqrt x \) will have the domain of negative integers also

OpenStudy (lgbasallote):

you lost me...

ganeshie8 (ganeshie8):

there is no inverse

OpenStudy (lgbasallote):

why so

OpenStudy (anonymous):

there is no inverse...

ganeshie8 (ganeshie8):

y = x^2 doesnt pass horizontal line test, in the domain of integers

hartnn (hartnn):

i^2 = -1 so if we allow x=i x^2 is negative then range of f(x) can be negative integers, so domain of f^-1(x) can be all integers....

OpenStudy (anonymous):

@ganeshie8 , Neither do inverse trig functions, but we can still find an inverse on an interval..

OpenStudy (anonymous):

the domain of f(x) will be the range of f^-1(x) if it exists....f^-1(x) can be sqrt(x) but its range should be positive integers but not all integers....complex plays no role here..

OpenStudy (lgbasallote):

makes sense